OpenStudy (anonymous):

I am working on problem set 4, 2C-4, and I am not understanding where they are converting x=18 (the critical point) to be 3/2 ft.. Is there something I am not seeing in this min/max problem? Thank you.

OpenStudy (phi):

the constraint is in inches: L+4w <= 108 inches solve to find L= 36 inches and w= 18 inches but they want the answer in cubic feet. 12 inches= 1 foot so L= 36 inches * 1 foot/12 inches = 3 ft w = 18 inches * 1 ft/12 in = 18/12 ft= 3/2 = 1.5 ft volume = (3/2)^2 * 3 ft^3 = 27/4 ft^3

OpenStudy (anonymous):

Ah, okay. That's right. Thank you!